OpenStudy (anonymous):

one batch of cookies are 2 cups of flour. How much flour do you need for 3/4 of a batch of cookies?

OpenStudy (anonymous):

4/3=2/f 4f=3*2 f=6/4 f=3/2

OpenStudy (anonymous):

1 bath of cookies need 3/2 flour

OpenStudy (anonymous):

4/3=2/f time and flip , I don't know how to wrtie faction ,